# DPZV: Elevating the Tradeoff between Privacy and Utility in Zeroth-Order Vertical Federated Learning

### v2

@@ Abstract
Vertical Federated Learning (VFL) enables collaborative training with feature-partitioned data, yet remains vulnerable to privacy leakage through gradient transmissions. Standard differential privacy (DP) techniques like DP-SGD are difficult to apply in this setting due to VFL’s distributed nature and the high variance incurred by vector-valued noise. On the other hand, zeroth-order (ZO) optimization techniques can avoid explicit gradient exposure, but lack formal privacy guarantees. In this work, we propose DPZV, the first ZO optimization framework for VFL that achieves tunable DP with performance guarantees. DPZV overcomes the aforementioned limitations by injecting low-variance scalar noise at the server, enabling controllable privacy with reduced memory overhead. We conduct a comprehensive theoretical analysis showing that DPZV matches the convergence rate of first order (FO) optimization methods while satisfying formal $(\epsilon,\delta)$-DP guarantees. Experiments on image and language benchmarks demonstrate that DPZV outperforms several baselines in terms of achieved accuracy under a wide range of privacy constraints ($\epsilon\leq 10$), thereby elevating the privacy-utility tradeoff in VFL.
1 Introduction
Vertical Federated Learning (VFL) has emerged as a compelling paradigm for training models on data that are partitioned feature-wise among distinct parties [15, 4, 3]. It seeks to preserve privacy by maintaining separate submodels locally and exchanging only intermediate representations or gradients, without exposing their raw data. Despite its privacy-aware architecture, recent studies have demonstrated that VFL remains vulnerable to feature leakage [19, 36] and label leakage [13, 41], where adversaries can infer sensitive input attribute or labels from the exchanged model updates or intermediate gradients. These findings highlight a critical need for integrating stronger, principled privacy-preserving mechanisms into the VFL framework.
One direction for mitigating privacy risks in VFL is through replacing SGD-based updates with zeroth-order (ZO) optimization[28]. By avoiding the transmission of intermediate gradients, ZO naturally reduces the risk of label leakage[39]. However, the privacy level of ZO-based methods are not directly adjustable, which constrains their applicability in settings that demand adaptive privacy guarantees aligned with task-specific requirements. Moreover, ZO remains susceptible to feature leakage, since sensitive information may still be inferred from local model parameters.
Importantly, current ZO-based VFL algorithms fail to provide mathematically rigorous differential privacy (DP) [10] guarantees, unlike their SGD counterparts. The DP mechanism is desirable as it allows for adjustable privacy levels under varying privacy budgets, enabling flexible trade-offs between model performance and privacy[11]. It also offers protection against a broad class of inference attacks  [40, 31, 22]. As a result, in this work, we ask the following research question:
How can we incorporate a controllable differential privacy mechanism into zeroth-order VFL which maintains convergence guarantees while elevating the privacy-utility tradeoff?
Key Challenges. Integrating differential privacy into ZO-based VFL poses several fundamental challenges. In particular, existing strategies for achieving DP within the VFL framework typically involve injecting vector-valued noise into either (i) forward embeddings or (ii) backward gradients[4, 35]. While approach (i) can provide formal privacy guarantees, it requires adding noise with the same shape as the transmitted embeddings, which are typically high-dimensional. This high-dimensional noise introduces substantial variance amplification[1]. Moreover, ZO methods inherently produce noisy gradient estimates[14]: when combined with the additional variance introduced by DP noise, the compounded effect can lead to substantial performance degradation. This situation can be further exacerbated when the privacy budget is tight. On the other hand, approach (ii) is incompatible with ZO methods, as ZO does not involve transmitting backward gradients[12].
Contributions. To address the above challenges, we propose Differentially Private Zeroth-Order VFL (DPZV), a novel ZO-based VFL framework that achieves a controllable differential privacy level with convergence guarantees. By injecting scalar-valued noise rather than high-dimensional vector noise, DPZV mitigates variance amplification, thereby enhancing the tradeoff between privacy and model performance in VFL. Our main contributions can be summarized as follows:
  • DPZV is the first ZO optimization framework for vertical federated learning (VFL) that enables tunable differential privacy. By injecting calibrated scalar-valued noise instead of high-dimensional vector noise, DPZV provides privacy guarantees while preserving model accuracy. Moreover, the use of ZO optimization substantially reduces memory overhead, making DPZV particularly suitable for privacy-critical and resource-constrained environments (Sec. 4).
@@ 1 Introduction
  • We provide a rigorous theoretical analysis establishing both convergence and privacy guarantees for DPZV. Specifically, we show that DPZV achieves a convergence rate on the same order as first-order DP-SGD, despite using only ZO estimators. This result indicates that our DPZV framework attains privacy without sacrificing training efficiency. Furthermore, we prove that our method satisfies $(\epsilon,\delta)$-DP, confirming its ability to provide an adjustable privacy control mechanism (Sec. 5 and 6).
@@
  • Through experiments on image and language benchmarks, we demonstrate that DPZV obtains robust convergence performance under strict privacy budgets. While other DP baselines experience a steep performance degradation as the privacy level increases, DPZV consistently maintains high accuracy across all tasks, underscoring an elevated tradeoff between privacy and utility (Sec. 7).
2 Related Work
Vertical Federated Learning. Vertical Federated Learning (VFL) enables collaborative training across organizations with vertically partitioned features. Early VFL frameworks focused on simple client-side models such as logistic regression and linear models [15], where clients computed local gradients for global aggregation. These methods prioritized simplicity, but lacked expressiveness for complex tasks. To address this limitation, larger client-side models like deep neural networks (DNNs) were adopted [4, 3, 35]. Studies have demonstrated that these larger client-side models can be more effective than classical methods [23]. However, larger models require higher processing capability and VRAM costs, which calls for computation and memory efficient methods.
In an effort to reduce communication cost, several communication-efficient VFL methods have been proposed. One popular mechanism for reducing communication overhead is by allowing for multiple local updates in-between aggregations. In this respect, [24] introduced FedBCD, a federated stochastic block coordinate descent method, which allows clients to perform multiple gradient iterations before synchronization. Similarly, Flex-VFL [3] proposed a flexible strategy offering varying local update counts per party, constrained by a communication timeout. VIMADMM [35] adopted an ADMM-based approach to enable multiple local updates in VFL. All these methods effectively reduce the number of round-trip communications in each iteration. On the other hand, Asynchronous VFL methods (e.g., FDML [17], VAFL [4]) decouple coordination, allowing clients to update models independently. While asynchronous methods risk staleness, they significantly improve scalability in cross-silo deployments. However, in first-order methods, the backward pass through neural networks typically imposes communication overhead comparable to the forward pass. In contrast, our ZO-based approach significantly reduces the cost associated with backward propagation.
Privacy guarantees are critical for VFL adoption. Some VFL architectures use crypto-based privacy-preserving techniques such as Homomorphic Encryption (HE) [6], but lack formal assurances, while DP provides rigorous mathematical protection. Key DP-based methods include VAFL [4], which injects Gaussian noise into client embeddings during forward propagation to achieve Gaussian DP, and VIMADMM [35], which perturbs linear model parameters with bounded sensitivity, ensuring DP guarantees for convex settings. Our work overcomes challenges in developing such methods for the ZO setting.
Zeroth-Order Optimization. Recent research has explored Zeroth-Order (ZO) optimization within VFL to accommodate resource-constrained clients with non-differentiable models and to reduce gradient leakage. Early work like ZOO-VFL [39] adopted a naive ZO approach throughout VFL training but provided no DP guarantees. VFL-CZOFO [33] introduced a cascaded hybrid optimization method, combining zeroth-order and first-order updates, which leveraged intrinsic noise from ZO for limited privacy. However, its DP level was not adjustable, often resulting in insufficient protection.
More recently, MeZO [26] proposed a memory-efficient ZO algorithm. Building upon these ideas, DPZero [38] and DPZO [32] introduced private ZO variants offering baseline privacy features. However, these methods are designed for centralized settings and cannot be directly extended to the VFL paradigm. In this work, we incorporate the memory-efficient design of MeZO into a VFL framework. Extending beyond previous efforts to combine ZO optimization with VFL, we further integrate a controllable differential privacy mechanism, achieving an elevated trade-off between privacy and model performance.
@@ 2 Related Work
Figure 1: Overview of the training procedure in DPZV. Each client perturbs its local model parameters in two random directions to generate a pair of embeddings, which are then transmitted to the server. The server computes the corresponding function evaluations and applies an elementwise difference to approximate the zeroth-order (ZO) gradient. To ensure differential privacy, scalar-valued Gaussian noise is injected into the aggregated ZO estimate. Unlike traditional vector-valued noise in standard DP algorithms, scalar noise is significantly smaller in norm, thereby preserving model utility even under stringent privacy budgets.
3 Background and Motivation
@@
System Model for VFL. We consider a VFL framework with one server and $M$ clients. In VFL, data is vertically partitioned between clients, with each client holding different features of the data. Suppose we have a dataset ${\mathcal{D}}$ with $D$ samples: ${\mathcal{D}}=\{\xi_{i}|i=1,2,\dots,D\}$. Each data sample $\xi_{i}$ can be partitioned into $M$ portions distributed throughout all clients, where the data sample on machine $m$ with ID $i$ is denoted as $\xi_{m,i}$, hence $\xi_{i}=[\xi_{1,i},\xi_{2,i},\ldots,\xi_{M,i}]^{\top}$. The server is numbered as machine 00 and holds the label data ${\bm{Y}}=\{y_{i}|i=1,2,\dots,D\}$.
@@
The clients and the server aim to jointly train a machine learning model parameterized by ${\bm{w}}$. The global model comprises local models on each party parameterized by $w_{m}$, with $m=0,1,\dots,M$ being the ID of the machine. To protect privacy, clients do not communicate with each other regarding data or model parameters. Instead, all clients communicate directly with the server regarding their local model’s outputs, which we term as local embeddings. If we denote the local embedding of client $m$ as $h_{m,i}:=h(w_{m};\xi_{m,i})$, the objective of the VFL framework can be seen as minimizing the following function:
$\textstyle F({\bm{w}};{\mathcal{D}},{\bm{Y}}):=\frac{1}{D}\sum_{i\in[D]}{\mathcal{L}}(w_{0},h_{1,i},h_{2,i},\dots,h_{M,i};y_{i})$ (1)
@@
$\textstyle f({\bm{w}};\xi_{i}):={\mathcal{L}}(w_{0},h_{1,i},h_{2,i},\dots,h_{M,i};y_{i})$ (2)
Memory-Efficient Zeroth-Order Optimizer. We introduce the two-point gradient estimator [30], which will serve as our zeroth-order gradient estimator throughout this paper. Let $u$ be uniformly sampled from the Euclidean sphere $\sqrt{d}{\mathbb{S}}^{d-1}$ and $\lambda>0$ be the smoothing factor. For a single datum $\xi$ sampled from the whole dataset ${\mathcal{D}}$, the two-point gradient estimator is defined as
$\textstyle g(x;\xi)=\frac{f(x+\lambda u;\xi)-f(x-\lambda u;\xi)}{2\lambda}u$ (3)
To further reduce the memory overhead on client, we adopt the MeZO methodology [26] for our ZO Optimization, which requires only the same memory as the model itself for training. We investigate the memory reduction in Appendix  F.
Differential Privacy and DP-SGD. Differential Privacy (DP) provides a principled framework for protecting individual data in statistical analysis and machine learning. Formally, we adopt the standard $(\epsilon,\delta)$-DP definition:
Definition 3.1 ($(\epsilon,\delta)$-Differential Privacy).

### v3

# Communication-Efficient and Differentially Private Vertical Federated Learning with Zeroth-Order Optimization

@@ Abstract
Vertical Federated Learning (VFL) enables collaborative model training across feature-partitioned devices, yet its reliance on device-server information exchange introduces significant communication overhead and privacy risks. Downlink communication from the server to devices in VFL exposes gradient-related signals of the global loss that can be leveraged in inference attacks. Existing privacy-preserving VFL approaches that inject differential privacy (DP) noise on the downlink have the natural repercussion of degraded gradient quality, slowed convergence, and excessive communication rounds. In this work, we propose DPZV, a communication-efficient and differentially private ZO-VFL framework with tunable privacy guarantees. Based on zeroth-order (ZO) optimization, DPZV injects calibrated scalar-valued DP noise on the downlink, significantly redu cing variance amplification while providing equivalent protection against targeted inference attacks. Through rigorous theoretical analysis, we establish convergence guarantees comparable to first-order DP-SGD, despite relying solely on ZO estimators, and prove that DPZV satisfies $(\epsilon,\delta)$-DP. Extensive experiments demonstrate that DPZV consistently achieves a superior privacy–utility tradeoff and requires fewer communication rounds than existing DP-VFL baselines under strict privacy constraints ($\epsilon\leq 10$).
I Introduction
Vertical Federated Learning (VFL) has emerged as a compelling paradigm for distributed training over edge and wireless systems where devices hold complementary features for the same data samples [hardy2017private, chen2020vafl, castiglia2023flexible]. Such vertical feature partitioning naturally arises in many applications. For example, in wireless sensor networks, spatially-distributed nodes may each collect local environmental measurements treated as features in a learning task [wang2018edge, shi2016edge, park2019wireless, guo2022nonparametric]. In other distributed setups, devices may only be capable of storing a portion of the features due to strict memory budgets or security concerns [yu2020low, liu2024vertical]. In these settings, VFL enables flexible model deployment by allowing learning to be distributed across devices with only a server/fusion center storing and processing the full model.
In most VFL systems, different devices holding partial features exchange intermediate gradients or embeddings with the server for aggregation [liu2020backdoor, chen2020vafl, weng2020practical]. This is in contrast to conventional (horizontal) FL [mcmahan2017communication, wang2018edge, 10.1109/MWC.011.2000501, 9276464], where each device holds a full feature vector for a subset of the samples, and updates are exchanged on the full model. The result is a different communication pattern in VFL, where information exchange typically occurs with respect to smaller submodel partitions rather than through full model updates. Still, communication efficiency is a pressing issue in VFL, influenced heavily by the feature dimensionality and training procedure [castiglia2023less, liu2024vertical].
Beyond efficiency concerns, VFL’s reliance on transmitting intermediate results introduces privacy concerns. Recent studies have shown that adversaries can exploit communicated information to infer sensitive attributes (feature inference attacks) [jin2021cafe, ye2024feature] or even recover labels (label inference attacks) [fu2022label, zou2022defending] and leverage them in e.g., backdoor attacks [lee2025cooperative]. Under the commonly adopted honest-but-curious threat model, these risks are particularly pronounced on the downlink, where gradient-related signals directly encode label-dependent information. Therefore, protecting downlink communication is essential for safeguarding privacy in VFL.
Unfortunately, existing mitigation measures for these privacy concerns often harm model performance and further exacerbate the communication burden of VFL. Similar to conventional FL, VFL methods often protect gradients by injecting calibrated noise vectors using differential privacy (DP) mechanisms [chen2020vafl, wang2024unified, xie2024improving, xu2019hybridalpha]. However, such noise inevitably degrades gradient quality, leading to slower convergence and thus requiring more communication rounds to reach a target accuracy. Consequently, privacy protection and communication efficiency are tightly coupled in DP-enhanced VFL (DP-VFL): stronger privacy requirements for a fixed target model quality typically translate into heavier communication costs.
I-A Zeroth-Order VFL
Zero-order (ZO) optimizers [nesterov2017random, fang2022communication] offer a compelling foundation for building private and efficient solutions to VFL [zhang2021desirable, wang2024unified]. Migrating from first-order gradient descent-based methods to zero-order VFL (ZO-VFL) is motivated by two key observations. First, ZO methods eliminate explicit gradient transmission in the backward phase, substantially reducing information leakage and providing a stronger baseline defense against label inference attacks [zhang2021desirable]. Second, ZO enables optimization using only function evaluations. As a result, the backward communication in ZO-VFL consists of scalar-valued feedback rather than high-dimensional gradients. This significantly limits the dimensionality of information transmitted and potentially revealed per downlink interaction.
However, ZO-VFL alone does not fully resolve privacy concerns. Despite the absence of explicit gradients, malicious devices can still approximate gradients from perturbation-based scalar feedback, leaving ZO-VFL vulnerable to inference attacks. Moreover, while recent work [wang2024unified] has quantified an inherent differential privacy provided by ZO methods, these levels are typically insufficient when strong DP guarantees are needed [gupta2024inherent]. Importantly, these DP levels are also not adjustable, limiting their practicality in real-world deployments that require controllable privacy guarantees.
These limitations lead to the central question of this work:
  • How can VFL obtain communication efficiency advantages as in ZO-VFL while preserving controllable privacy levels as in DP-VFL?
To answer this question, we observe that achieving communication efficiency in DP-VFL is fundamentally tied to obtaining a favorable privacy–accuracy tradeoff, since improved convergence behavior directly translates into fewer communication rounds needed to reach a target accuracy. However, accomplishing this goal faces two key challenges. First, existing approaches for enforcing differential privacy in VFL commonly inject noise into the forward embeddings [chen2020vafl, xie2024improving], which are typically high-dimensional. As a result, the injected noise induces substantial variance amplification and significantly slows convergence, particularly under tight privacy budgets. Second, ZO optimization methods rely on inherently noisy gradient estimators and are generally known to converge more slowly than first-order methods [ghadimi2013stochastic]. When combined with additional DP noise, this effect can be further amplified, potentially leading to severe performance degradation and an increased number of communication rounds.
I-B Summary of Contributions
In this work, we propose DPZV, a VFL methodology which achieves communication efficiency and controllable differential privacy concurrently. Instead of perturbing the forward embeddings, our method injects scalar-valued DP noise into the backward information. Intuitively, injecting scalar noise preserves the directional structure of the ZO gradient estimator and introduces far less distortion than vector-valued perturbations, enabling more accurate updates and faster convergence under the same privacy budget. We show in Sec. VI that this design achieves the same level of protection against targeted label and feature inference attacks as forward perturbations, while significantly reducing noise dimensionality.
Through both theoretical analysis and extensive experiments, we demonstrate that our approach provides faster convergence than existing methods under a fixed privacy budget, incurring the lowest total communication cost to reach a target accuracy. More specifically, our contributions are as follows:
  • DPZV is the first VFL methodology that enables tunable differential privacy while maintaining communication efficiency. Unlike existing DP-VFL methods that inject high-dimensional noise into forward embeddings, DPZV perturbs only the scalar-valued backward information. This design preserves privacy guarantees while substantially reducing variance in the gradient estimation, leading to faster convergence and fewer required communication rounds. By eliminating gradient transmission, DPZV is particularly well suited for privacy-critical and communication-constrained deployments (Sec. IV).
@@ I-B Summary of Contributions
  • We provide a rigorous theoretical analysis establishing both convergence and privacy guarantees for DPZV. Specifically, we show that DPZV achieves a convergence rate to a stationary point on the same order as first-order DP methods, despite using less precise ZO estimators. This result demonstrates that the conventional gap in convergence speed between ZO and first-order methods diminishes in the DP-VFL setting. Furthermore, we prove that our method satisfies $(\epsilon,\delta)$-DP, confirming its ability to provide an adjustable privacy control mechanism while retaining communication efficiency (Sec. V, VI).
@@
  • Through experiments on four benchmark datasets, we verify that DPZV obtains robust convergence performance under strict privacy budgets. In particular, while the baseline algorithms experience a steep performance degradation and/or increase in communication rounds as the privacy level increases, we find that DPZV consistently maintains faster convergence to a target accuracy across all tasks, underscoring an elevated privacy-utility-cost tradeoff. In addition to a lower total communication cost, we find that DPZV reduces the required memory overhead for training large models (Sec. VII).
@@
Figure 1: Overview of the training procedure in DPZV. Each device perturbs its local model parameters in two random directions to generate a pair of embeddings, which are then transmitted to the server. The server computes the corresponding function evaluations and applies an elementwise difference to approximate the zeroth-order (ZO) gradient. To ensure differential privacy, scalar-valued Gaussian noise is injected into the aggregated ZO estimate. Unlike traditional vector-valued noise in standard DP algorithms, scalar noise is significantly smaller in norm, thereby preserving model utility even under stringent privacy budgets.
II Related Work
II-A Vertical Federated Learning
VFL enables collaborative training across entities (e.g., organizations or, in our case, wireless devices) with vertically partitioned features. Early VFL frameworks focused on simple device-side models such as logistic regression and linear models [hardy2017private]. These methods prioritized simplicity, but lacked expressiveness for complex tasks. To address this limitation, larger device-side models like deep neural networks (DNNs) were adopted [chen2020vafl, castiglia2023flexible, xie2024improving].
A key challenge in VFL is the communication overhead incurred during training. One popular mechanism for reducing communication overhead is by allowing for multiple local updates in-between aggregations. In this respect, [liu2022fedbcd] introduced FedBCD, which allows devices to perform multiple gradient iterations in VFL before synchronization. Similarly, Flex-VFL [castiglia2023flexible] proposed a flexible strategy offering varying local update counts per party, constrained by a communication timeout. VIMADMM [xie2024improving] adopted an ADMM-based approach to enable multiple local updates in VFL. On the other hand, Asynchronous VFL methods (e.g., FDML [hu2019fdml], VAFL [chen2020vafl]) decouple coordination, allowing devices to update models independently, thus improving scalability. However, in first-order methods, the backward pass through neural networks typically imposes communication overhead, while our ZO-based approach significantly reduces the cost associated with backward propagation.
Privacy guarantees are another critical challenge for VFL adoption. Some VFL architectures use crypto-based privacy-preserving techniques such as Homomorphic Encryption (HE) [cheng2021secureboost], but lack formal assurances. In contrast, DP provides rigorous mathematical protection. Key DP-based methods include VAFL [chen2020vafl], which injects Gaussian noise into device embeddings during forward propagation to achieve Gaussian DP, and VIMADMM [xie2024improving], which perturbs linear model parameters with bounded sensitivity, ensuring DP guarantees for convex settings. Our work overcomes challenges in developing such methods for the ZO setting.
II-B Zeroth-Order Optimization
Recent research has explored ZO optimization within VFL to accommodate resource-constrained devices with non‐differentiable models and to reduce gradient leakage. Early work like ZOO-VFL [zhang2021desirable] adopted a naive ZO approach throughout VFL training but provided no DP guarantees. VFL-CZOFO [wang2024unified] introduced a cascaded hybrid optimization method, combining zeroth-order and first-order updates, which leveraged intrinsic noise from ZO for limited privacy. However, its DP level was not adjustable, resulting in insufficient protection.
More recently, MeZO [malladi2023fine] proposed a memory-efficient ZO algorithm. Building upon these ideas, DPZero [zhang2024dpzero] and DPZO [tang2024private] introduced private ZO variants offering baseline privacy features. However, these methods are designed for centralized settings and cannot be directly extended to the VFL paradigm. Extending beyond previous efforts to combine ZO optimization with VFL, we further integrate a controllable differential privacy mechanism, achieving an elevated trade-off between privacy and model performance.
III Preliminaries
III-A System Model for VFL
@@
We consider a VFL framework with one server and $M$ devices. In VFL, data is vertically partitioned between devices, with each device holding different features of the data. Suppose we have a dataset $\mathcal{D}$} with $D$ samples: ${\mathcal{D}}=\{\xi_{i}|i=1,2,\dots,D\}$. Each data sample $\xi_{i}$ can be partitioned into $M$ portions distributed throughout all devices, where the data sample on machine $m$ with ID $i$ is denoted as $\xi_{m,i}$, hence $\xi_{i}=[\xi_{1,i},\xi_{2,i},\ldots,\xi_{M,i}]^{\top}$. The server is numbered as machine 0 and holds the label data ${\bm{Y}}=\{y_{i}|i=1,2,\dots,D\}$.
@@
The devices and the server aim to jointly train a machine learning model parameterized by $\bm{w}$}. The global model comprises local models on each party parameterized by $w_{m}$, with $m=0,1,\dots,M$ being the ID of the machine. To protect privacy, devices do not communicate with each other regarding data or model parameters. Instead, all devices communicate directly with the server regarding their local model’s outputs, which we term as local embeddings. If we denote the local embedding of device $m$ as $h_{m,i}:=h(w_{m};\xi_{m,i})$, the objective of the VFL framework can be seen as minimizing the following function:
$F({\bm{w}};{\mathcal{D}},{\bm{Y}}):=\frac{1}{D}\sum_{i\in[D]}{\mathcal{L}}(w_{0},h_{1,i},h_{2,i},\dots,h_{M,i};y_{i})$ (1)
@@
$f({\bm{w}};\xi_{i}):={\mathcal{L}}(w_{0},h_{1,i},h_{2,i},\dots,h_{M,i};y_{i})$ (2)
Memory-Efficient Zeroth-Order Optimizer. We introduce the two-point gradient estimator [shamir2017optimal], which will serve as our zeroth-order gradient estimator throughout this paper. Let $u$ be uniformly sampled from the Euclidean sphere $\sqrt{d}{\mathbb{S}}^{d-1}$ and $\lambda>0$ be the smoothing factor. For a single datum $\xi$ sampled from the whole dataset $\mathcal{D}$}, the two-point gradient estimator is defined as
$g(x;\xi)=\frac{f(x+\lambda u;\xi)-f(x-\lambda u;\xi)}{2\lambda}u$ (3)
To further reduce the memory overhead on device, we adopt the MeZO methodology [malladi2023fine] for our ZO Optimization, which requires only the same memory as the model itself for training. We investigate the memory reduction of this estimator in Section VII.
III-B Differential Privacy and DP-SGD
DP provides a principled framework for protecting individual data in statistical analysis and machine learning. Formally, we adopt the standard $(\epsilon,\delta)$-DP definition:
Definition III.1 ($(\epsilon,\delta)$-Differential Privacy).